OpenStudy (anonymous):

does anyone know what 5-n=1 n= or (there are absolute value signs around the 5-n)

OpenStudy (anonymous):

\[|5-n|=1?\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

you solve for both of the following: \[5-n=1~~~~~~~~~~~~~~~~~~~~~~~5-n=-1\] which give you: \[n=4~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~n=6\]

OpenStudy (anonymous):

so your answer is \[n=~4 ~~or~~6\]

OpenStudy (anonymous):

thanks but how you know this

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

youre welcome

OpenStudy (anonymous):

|2-n|=3

OpenStudy (anonymous):

do the same: solve for both situations \[2-n=3~~~~~~~~~~~~~2-n=-3\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

would it be 5 and -5

OpenStudy (anonymous):

no.. \[2-n=3\]\[-n=3-2\]\[-n=1\]\[n=-1\] and \[2-n=-3\]\[-n=-3-2\]\[-n=-5\]\[n=5\] so the answers are -1 and 5

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

youre welcome :)